media%2F4e0%2F4e0dcc45-63e3-4f6f-b341-7c
Answer #1

(10)

(a) Balanced equation

4NH3 + 5O2 -------------> 4NO + 6H2O

(b) Moles of ammonia used = 60 * 1000 / 17 = 3529.4 mol

Moles of Oxygen used = 110 * 1000 / 32 = 3437.5 mol

According to balanced equation, 4 mole of ammonia requires 5 moles of oxygen.

Hence, oxygen is limiting reagent. (3437.5 < 3529.4)

From the balanced equation,

5 moles of oxygen can form 4 moles of NO and 6 moles of H2O

Then, 3437.5 moles of O2 can from 3437.5 * 4 / 5 moles of NO and 3437.5*6/5 moles of H2O

SO, moles of NO formed = 2750 mol

Mass of NO formed = 2750 * 30 /1000 = 82.5 kg.

Similarly, moles of H2O fromed = 4125 mol

Mass of H2O formed = 4125 * 18 / 1000 = 74.25 kg.